OpenStudy (anonymous):

Trevor solved the system of equations below. What mistake did he make in his work? 2x + y = 5 x − 2y = 10 y = 5 − 2x x − 2(5 − 2x) = 10 x − 10 + 4x = 10 5x − 10 = 10 5x = 0 x = 0 2(0) + y = 5 y = 5 He should have substituted 5 + 2x He combined like terms incorrectly, it should have been 4x instead of 5x He subtracted 10 from the right side instead of adding 10 to the right side He made no mistake

OpenStudy (anonymous):

He made no mistakes. Follow through his work and do it yourself and you will find that those answer are correct. You could also plug the values in for x and y and see if they check out.

OpenStudy (anonymous):

ok thank you

OpenStudy (phi):

5x − 10 = 10 5x = 0 Do you see anything wrong with this step? I do.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i do see that i didnt see that before

OpenStudy (phi):

what happens if you add +10 to both sides? Like this: 5x − 10 +10 = 10+10

OpenStudy (anonymous):

Sorry, he is right. I should have been wearing my glasses.

OpenStudy (phi):

do you see that -10 +10 is 0 ? 5x − 10 +10 = 10+10 becomes 5x = 10+10

OpenStudy (phi):

and 10+10 is 20 so you get 5x=20 which does not match what Tevor did in the problem

OpenStudy (anonymous):

srrry was afk so its C?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thank you guys